State Boyle's law and Charles' law and show how they may be combined to give the equation of state of an ideal gas.
1
Expert's answer
2021-02-08T18:52:34-0500

Boyle’s law:

Boyle’s law states the relation between volume and pressure at constant temperature and mass.

"p \u221d 1\/V"

if a fixed mass of gas undergoes an expansion at constant temperature then the final volume and pressure shall be pand V2. The initial volume and initial pressure here is p1 and V1 then according to Boyle’s law: p1×V1 = p2×V2 = constant (k1)


"\frac{p_1}{p_2} = \frac{V_2}{V_1}"


Charles' law :

Charles' law at constant pressure and for constant mass, the volume of a gas is directly proportional to the temperature.


"\frac{V}{T} = constant" or


"\frac{V1 } {T1} = \frac{V2 } { T2}"